determine the total charge (a) on line 0 < x < 5 m if pl = \2x2 mc/m


1
Expert's answer
2021-05-24T18:38:02-0400

Q"_l=\int\times dx"

"=\int P_ldx"

"=\int_0^5 12x^2dx\times10^{-3}"

"=12\frac{x^3}{3}\int_0^5\times10^{-3}"

"=12\times\frac{(5)^3}{3} \times10^{-3}"

"=500\times10^{-3}"

Q"_l=500mC"